OpenStudy (anonymous):

Jamal works in a city and sometimes takes a taxi to work. The taxicabs charge $1.50 for the first 1/5 mile and $0.25 for each additional 1/5 mile. Jamal only has $3.75 in his pocket. What is the maximum distance he can travel by taxi if he does not tip the driver?

OpenStudy (anonymous):

thanks man i appreciate it

OpenStudy (anonymous):

how many miles to his work

OpenStudy (anonymous):

it didn't say sorry

OpenStudy (anonymous):

ok that weird

OpenStudy (anonymous):

i think all you need to know is how much money he has in his wallet

OpenStudy (anonymous):

@Ashleyisakitty please help me

OpenStudy (anonymous):

yea true

OpenStudy (anonymous):

2 miles i could be wrong but that's what i got

OpenStudy (anonymous):

could you show the work please

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

3.75-1.50= 2.25 then u need to do is subtract 2.25-0.25=2 which equals 2 that's how i got my answer

OpenStudy (anonymous):

ok thanks man

OpenStudy (anonymous):

ur welcome

OpenStudy (triciaal):

the actual question is how many miles can he travel and spend less or = $3.75? "i think all you need to know is how much money he has in his wallet" you are told he has $3.75 1.50 + 0.25(x-1/5) </= 3.75 where x is the number of miles 1.50 +0.25x - 0.25(1/5) < 3.75 0.25x < 3.75 - 1.50 + 0.05 x < 2.3/0.25 x < 9.2